OpenStudy (banjoman):

cash register contains only five dollar and ten dollar bills. It contains twice as many five's as ten's and the total amount of money in the cash register is 700 dollars. How many ten's are in the cash register? Will medal

OpenStudy (brooke..help00):

2X*5+X*10=680 Multiply 2X by 5 to get 10X. 10X+X*10=680 Multiply X by 10 to get 10X. 10X+10X=680 Since 10X and 10X are like terms, add 10X to 10X to get 20X. 20X=680 Divide each term in the equation by 20. (20X)/(20)=(680)/(20) Simplify the left-hand side of the equation by canceling the common factors. X=(680)/(20) Simplify the right-hand side of the equation by simplifying each term. X=34 tens 68 fives Prove: 34*10= 340 68*5= 340 340+340=680

OpenStudy (sissyedgar):

There would be 35 tens because... There is a total of $700. 700 / 2 = 350 350 / 10 = 35 We already know that 5 * 2 = 10... So whatever the amount of $10 bills is, you just times it by 2 and you have the amount of fives. 35 * 2 = 70.